OpenStudy (anonymous):

Use the properties of exponents to simplify the expression, Write with positive exponents. Assume that all variables request positive real numbers. (y^1/2)/(y^1/4)

jhonyy9 (jhonyy9):

us property of (x^a)/x^b) = x^(a-b)

jhonyy9 (jhonyy9):

- so than will be y^(1/2 -1/4) =y^((2-1)/4) =y^(1/4)

OpenStudy (anonymous):

Thank-you